OpenStudy (anonymous):

10 is multiplied by the sum of 3 and 2. The product is divided by 2, and then 5 is added.

OpenStudy (anonymous):

The answer is 30. 3+2=5. 10*5=50. 50/2=25. 25+5=30.

OpenStudy (anonymous):

30 because 10 times 3+2 is 50 divided by 2 then adding 5